OpenStudy (anonymous):

Write the equation for a circle centered at the origin with a radius of 14.

OpenStudy (anonymous):

(x - x14)2 + (y - y14)2 = r2

OpenStudy (anonymous):

(x - 0)2 + (y - 0)2 = 142

OpenStudy (anonymous):

x2 + y2 = 196

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Now give me the medal

OpenStudy (anonymous):

Please

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (anonymous):

Wait, are you sure that's right? How did you get 142?

OpenStudy (anonymous):

I got 196

OpenStudy (aravindg):

14^2

OpenStudy (aravindg):

196^

OpenStudy (anonymous):

Radius is 14 and it says r2 So 14*2 28

OpenStudy (aravindg):

the general equation fo a circle centred at origin is x^+y^2=r^2

OpenStudy (aravindg):

Its r^2 not r*2

OpenStudy (anonymous):

666 fans ima make it 667

OpenStudy (aravindg):

r^r=r*r

OpenStudy (anonymous):

OH. OKAY. Sorry, I was confused.

OpenStudy (aravindg):

:) np